y=-x+16
y-x=2
I have to use substution method.
I got x=-5 and y=-3, but i doubt it is right

Answer by unlockmath(1688)

Hello,
Let's march through this and see what we get.
y=-x+16
y-x=2
Let's substitute the y in the first equation into the second one to look like:
-x+16-x=2 Now combine like terms and subtract 16 from both sides to look like:
-2x=-14 Divide -2 into both sides to give us:
x=7
Now we can plug 7 into either equation to find x.
y=-7+16 Combine terms to result in:
y=9
check these answers by plugging 7 and 23 into the original equations.
RJ
Check out a new book I wrote at:
www.math-unlock.com